# The physical basis of teleological evolutions

### v4

@@ 1 Introduction
The notion of teleological evolution is that of an evolution toward a goal with an attractor in the same goal it will produce in the future. It is, of course, a puzzling notion since it necessarily implies a causal loop. As is well known, it was often applied to the evolutions of the living before the scientific revolution. With the advent of the latter, it was dismissed because of its alleged lack of physical basis. That was, of course, under the notion of causal evolution that is dominant in science even today.
We take the opposite position in the present work. We will show that, in quantum mechanics, there is a well defined physical basis to the notion of teleological evolution.
In the first place, we show that quantum algorithms are teleological evolutions, indeed evolutions toward a goal (the solution) with an attractor in the goal they will produce in the future (the solution again).
Of course, seeing this requires a quantum description of quantum algorithms different from the ordinary one, which is naturally that of a causal process. The latter quantum description, with causality always going in the same time-direction, cannot describe causal loops.
We show that the ordinary quantum description of quantum algorithms, although complete from the mathematical standpoint, is physically incomplete and is completed by time-symmetrizing it. This leaves the ordinary quantum description mathematically unaltered but replaces its single causal process by a quantum superposition of causal loops. In each of them, it is as if the problem-solver knew in advance, before beginning her problem-solving action, one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution with fewer logical operations (computation steps). Of course, we are dealing with an evolution that responds to the notion of teleological evolution: it is indeed toward a goal (the solution) with an attractor in the goal it will produce in the future (the solution again).
This completion of the quantum description provides a unified quantitative explanation of all the optimal quantum speedups achieved in solving oracle problems. Given an oracle problem, the number of computation steps needed to solve it in an optimal quantum way is that needed to solve it in a optimal logical way benefiting of the advanced knowledge of any of the possible halves of the information about the solution. Of course, there is no such an explanation in current quantum computing science. Obviously, we have a causal quantum description for each quantum algorithm, but there is no known way of unifying them.
This part of the work is a further development of the retrocausal approach to the quantum computational speedup $\left[1-5\right]$.
Of course, both the approach in question and the part of the work discussed so far are unconventional and potentially controversial. The notion of causal loop is puzzling and mostly extraneous to today’s physics. However, we believe that this time we have given a very simple and easily checkable demonstration that the quantum computational speedup is physically due to quantum causal loops.
@@
Let us outline it. We consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem. If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case – by the optimal quantum search algorithm of Grover $\left[6\right]$ – she always solves the problem by opening just one drawer222 In the four drawers case, Grover algorithm yields the solution with certainty. In the following, by Grover algorithm we will always mean Long’s $\left[7\right]$ version of it, which does that always. . This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of viewing it.
@@
It should already be clear from the above that that of the quantum speedup is a fundamental discovery that changes a long standing way of viewing things333 The notion that computation is possible in the quantum world, with the related notion of quantum bit, was introduced by Finkelstein $[8]$ in 1969. In 1982, Feynman $\left[9\right]$ observed that the classical simulation of a quantum process can require an exponentially higher amount of physical resources. In 1985, Deutsch $\left[10\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup. .
We can go now to our demonstration, which is surprisingly simple. Of course, we must adopt a way of thinking open to retrocausality.
@@
The fact that Alice always solves the problem of locating the ball by opening just one drawer is logically equivalent to her knowing in advance, before beginning her problem-solving action, that the ball is in a given pair of drawers, in a mutually exclusive way one of the three pairs obtained by pairing the drawer with the ball with another drawer. In fact, if the ball is in the drawer opened, she has solved the problem. If not, it must be in the other drawer and she has solved the problem of locating the ball as well, having opened just one drawer.
@@
Then the question becomes: from where does Alice always get the information that the ball is in a given pair of drawers? Since, to her, the only possible source of it is the solution of the problem she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution. This is obviously a causal loop. Alice knows in advance one of the possible halves of the information about the solution she will produce and measure in the future and can use this knowledge to produce the solution by opening correspondingly fewer drawers. This, in the mutually exclusive $or$ of all the possible ways of taking half of the information about the solution.
We would argue that, with the above elementary logic, we have, so to speak, broken the sound barrier: the quantum speedup does logically imply causal loops. Then, one gos from logic to physics as follows.
Involving only the measurements of commuting observables (the problem-setting and the corresponding solution), quantum algorithms are submitted to the same (classical) logic. Therefore they must host the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of an optimal quantum speedup; of course, logical mutual exclusivity of possibilities, moving to the quantum world, becomes their quantum superposition. We have already seen that the quantum description of this quantum superposition is generated by time-symmetrizing the ordinary quantum description of the quantum algorithm. This time-symmetrization must therefore be performed because, as we will see, is the only way of generating the causal loops that must be hosted in the quantum algorithm.
Let us add a few comments to the above.
By ordinary quantum description of quantum algorithms, we do not mean their usual quantum description. The latter is limited to the unitary process that solves the problem, where the quantum speedup resides, and the final measurement of the solution. It is trivially incomplete since it lacks the initial measurement. Since we need to complete the quantum description, in the first place we should emend this trivial incompleteness. We do this by extending the usual quantum description to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings.
In the case of Grover algorithm, we obtain a unitary evolution between an initial and a final one-to-one correlated measurement outcomes encoding respectively the problem-setting (the number of the drawer with the ball selected by the initial measurement) and the corresponding solution of the problem. We have called it the ordinary quantum description of the quantum algorithm, ordinary since it is still a causal description that cannot describe causal loops. Note that any quantum algorithm that produces the solution with certainty can be put in this form. If the solution is not an invertible function of the problem-setting, it suffices that Alice, at the end, measures both the solution and the problem setting. This is a trivial but important step: of course, the usual quantum description of quantum algorithms would completely obscure the approach we are following.
Let us eventually note that we are dealing with the same kind of incompleteness of the ordinary quantum description that was ascribed by Einstein, Podolsky and Rosen $\left[11\right]$ to the quantum description of quantum nonlocality. Also this quantum description can be seen as a unitary evolution (the process of spatial separations of the two subsystems) between two one-to-one correlated measurement outcomes (those of measuring in the same basis the two respective maximally entangled observables) – we must think that this process goes from one to the other measurement outcome via the time the two subsystems were not spatially separate. Time symmetrized, it generates a quantum superposition of causal loops that play the role of the hidden variables envisaged by those authors to physically (locally) explain quantum non-locality $\left[5\right]$.
This ends the first part of the work, a review and further clarification of the retrocausal approach to the quantum speedup.
Of course, under the above explanation of their quantum speedup, quantum algorithms turn out to be teleological evolutions. The second part of the work is about the possibility that the quantum causal loops that stand at the basis of the teleological character of quantum algorithms are the missing physical basis of the teleological character of the evolutions of the living.
This immediately raises two problems.
One is explaining the presence of quantum features (the quantum speedup) in evolutions that can be the size of the Earth and have durations of billions of years. The obvious impossibility of scaling up the quantum features of the microscopic world to such levels points to the only possible solution: to scale them down from the largest and longest existing evolution there is, that of the universe. Of course, at the same time, we should adopt quantum cosmology, that is, the assumption that the universe – a perfectly isolated system since by definition there is nothing outside of it – evolves in a unitary quantum way.
The other is to find a formal analogy between quantum algorithms and, under quantum cosmology, natural evolutions. Obviously, it should justify the fact that also natural evolutions can occur with a quantum speedup and are consequently teleological in character.
Interestingly, the Anthropic Principle comes to our rescue. The Fine-tuned Universe version of it asserts that the unitary evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today, in particular (“sentient life” or “no life”). The reason for this correlation is that the slightest change of the value of the fundamental constants would produce a radically different universe most probably unable to host life.
Having adopted quantum cosmology, of course, also implies a quantum interpretation of the Anthropic Principle. Simply, the correlation between the value of the physical constants since the beginning of time and the value of life today becomes quantum and its building up is performed by the universe’s unitary quantum evolution.
This is what we were looking for. In fact, it is exactly what quantum algorithms do. They unitarily build up a quantum correlation between the problem-setting since the beginning of them and the solution of the problem at their end. Under this formal analogy, the evolution of the universe toward the value of life could be endowed with quantum speedup and would even in the case that the quantum search for the correlation in question is random, exactly as in random quantum search algorithms $\left[12\right]$.
Of course, being endowed with quantum speedup, under the quantum causal loops explanation of it, the evolution in question is teleological in character. In the universe we are in, what becomes teleological in character is the evolution of the universe toward sentient life.
Eventually, as we shall see, the evolution of life on Earth, in principle obtainable from that of the universe toward life by tracing out everything else, even if decohered, should inherit the quantum speed and thus the teleological character of the latter evolution. The same would consequently hold for any evolution of life on Earth, e. g. that of species.
The notion of a teleological evolution of life is of course puzzling. For example, the Darwinian theory of evolution would change as as follow: those random DNA mutations that, in the future, will survive natural selection would become more probable. With the past influenced by the future, this is a completely new way of viewing the physical world. No wonder the teleological notion has been rejected by science.
In conclusion, the work is highly unconventional, but its findings were obtained in a surprisingly simple and well-verifiable way. If it were correct, as we think, it would be a paradigm shift not only for the natural sciences, but also for the way we see physical reality.
2 Rationale
As the present work is about a fundamentally different way of viewing the role of causality in quantum (correlation) processes, before going to the formalizations we provide a Rationale that describes it conceptually.
This Rationale is divided into two parts corresponding to the two parts in which the work is divided.
The first provides the concept of the time-symmetric quantum causal loops, showing how they stand at the physical basis of the quantum computational speedup and quantum nonlocality. In particular, the time-symmetric quantum causal loop description of quantum algorithms highlights their teleological character, their being evolutions toward a goal (the solution of the problem) with an attractor in the goal they will produce in the future (the solution of the problem again).
The second addresses the question whether the time-symmetric quantum causal loops that stand at the basis of the teleological character of quantum algorithms could also be the missing physical basis of a teleological character of the natural evolutions for which the teleological notion was conceived.
2.1 The time-symmetric quantum causal loops – the concept
We go now to the first part of the work. We begin by showing, by the way very simply, that the sheer existence of the optimal quantum speedup delivered by Grover $\left[6\right]$ quantum search algorithm logically implies the mutually exclusive $or$ of well defined causal loops. Of course, we must ignore the ban on the use of the notion of retrocausality.
The fact that Alice (the problem-solver) always solves the problem of locating the ball by opening just one drawer logically implies that it is as if she always knew in advance, before beginning her problem-solving action, that the ball is in one of the three pairs of drawers formed by pairing the drawer with the ball with another drawer.
Indeed, in any of these three mutually exclusive cases, she opens either drawer; if the ball is in it, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well.
The question is then: from where does Alice always get (in a mutually exclusive way) this advanced knowledge of one of the possible halves of the information that specifies the solution? Since, to her, the only possible source of it is the solution she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution. This is, of course, a causal loop that allows her to produce the solution by opening just one drawer. There must be, in a mutually exclusive way, one causal loop for each of the three possible ways of knowing in advance one of the possible halves of the information that specifies the solution.
@@ 2.1 The time-symmetric quantum causal loops – the concept
We consider now $N$ drawers. Grover algorithm (also in Long’s version) yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case. Also now, this quantum speedup is logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is in one of the possible tuples of $\sqrt{N}$ drawers. She then solves the problem, in the best possible logical way, by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
We can therefore formulate the following advanced knowledge rule:
@@
The number of computation steps (drawer openings) needed to solve Grover’s problem in an optimal quantum way is that needed to solve it in an optimal logical way given the advanced knowledge of one of the possible halves of the information about the solution.
Of course, involving only the measurements of commuting observables (the problem-setting and the corresponding solution), Grover algorithm, as any other quantum algorithm, is submitted to the same (classical) logic. Its quantum description should therefore describe the quantum superposition of the causal loops in question (logical mutual exclusivity of possibilities becomes their quantum superposition in the quantum world).
However, the above immediately raises a problem. The current quantum description of quantum algorithms, as that of any quantum process, is causal in character. In other words, causality always goes in the same time direction. Obviously, it is a description that cannot describe causal loops; we tentatively ascribe this to its incompleteness.
Before going to this assumed incompleteness, of a fundamental kind, we should emend a trivial incompleteness of the current quantum description of quantum algorithms. It is limited to the unitary process of solving the problem, where the quantum speedup resides, and the final measurement of the solution. It is obviously incomplete because it lacks the initial measurement. Then, we extend it to the upstream process of setting the problem by measuring the problem-setting in a quantum superposition (or, indifferently, mixture) of all possible problem-settings. In the case of Grover algorithm, where the solution is an invertible function of the problem-setting (both are the number of the drawer with the ball), we obtain a unitary evolution (encoding the process of solving the problem) between an initial and a final one-to-one correlated measurement outcomes (encoding respectively the problem-setting selected by the initial measurement and the corresponding solution read by the final measurement). Of course, this unitary evolution never changes the problem-setting selected by the initial measurement.
By the way, seeing quantum algorithms as unitary evolutions between an initial and a final one-to-one correlated measurement outcomes is a key step of the present work. In view of going beyond Grover algorithm, we note that, if the solution is not an invertible function of the problem-setting, we can have one-to-one correlation between the initial and final measurement outcomes as well by requiring that, at the end, the problem-solver measures both the solution and the problem setting. The only constraint is that the quantum algorithm determines the solution with certainty.
We call the above, less incomplete, quantum description of the quantum algorithm its ordinary quantum description, ordinary because it is still a causal description that cannot describe causal loops. We should complete this ordinary quantum description in a way that leaves it mathematically unaltered but changes its causality structure – describes the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup. This is simply done time-symmetrizing it as follows.
Instead of ascribing the selection of the pair of one-to-one correlated measurement outcomes all to the initial measurement, as customary, we require that it time-symmetrically shares between the initial and final measurements, of course in a quantum superposition of all the possible ways. In each time-symmetrization instance, the initial measurement selects one of the possible halves of the information that specifies the sorted out pair of correlated measurement outcomes (it selects a quantum superposition of problem-settings instead of the customary single one). Its measurement outcome unitarily propagates forward in time until becoming the state immediately before the final measurement, a quantum superposition of tensor products each a problem-setting multiplying the corresponding solution. The final measurement completes the selection, of course by selecting a single pair problem-setting and corresponding solution. Its measurement outcome, obviously the same of the ordinary quantum description, time-symmetrically propagates backward in time, by the inverse of the time-forward unitary transformation, until replacing the previous initial measurement outcome.
This backward in time propagation, which inherits both selections, is an instance of the time-symmetrization of the ordinary quantum description of the quantum algorithm. Note that, mathematically, i.e. disregarding the time-direction of causality, it is the ordinary quantum description back again.
Also note that, together, the forward and backward in time propagations describe a causal loop that changes its own past (we call it time-symmetric since it is generated by a time-symmetrization procedure). As seen by the problem-solver, from whom the outcome of the initial measurement should always be concealed (would give her information about the solution – see Subsection 4.2.), it is one of the causal loops logically implied by the sheer existence of the optimal quantum speedup.
Eventually, we should take the quantum superposition of all the causal loops. Of course, the backward in time propagations of all time-symmetrization instances are identical with each other. Therefore, as required, this time-symmetrization leaves the unitary evolution between the two measurement outcomes mathematically unaltered. However, it changes the structure of causality along it. As seen by the problem-solver, the single causal process of the ordinary quantum description is replaced by the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of the quantum speedup. As already said, in each causal loop it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future; then, she can use this knowledge to produce the solution with fewer computation steps. If she exploits it entirely, the quantum speedup is optimal.
In equivalent terms, in each causal loop, the computational complexity of the problem to be solved reduces – this will become tangible in the formalizations of Section 4. The number of computation steps needed to solve the reduced complexity problem in an optimal quantum way is that needed to solve it in an optimal logical way. In still equivalent terms, the dimension of the Hilbert space in which the search for the solution takes place correspondingly reduces.
In view of what will follow, it is important to note that these results have been obtained independently of the distinctive character of the unitary evolution between the two correlated measurement outcomes (i.e. the actual quantum algorithm). It sufficed to know that it is a unitary evolution that one-to-one correlates two initially unrelated observables (the problem-setting observable and the solution observable) never changing the eigenvalue of the observable selected by the initial measurement.
In the present work, we will always know that there is such a unitary evolution. It is Grover algorithm in the present case, but it could be any other quantum algorithm that provides the solution with certainty – see also $\left[4\right]$. Under the assumption that between two one-to-one correlated measurement outcomes there can always be a unitary evolution of the kind in question, the advanced knowledge rule holds for any oracle problem.
Eventually, we note that the time-symmetrization in question, and the consequent time-symmetric quantum causal loops, are imposed by logic and quantum mechanics. Summarizing, logic says that the sheer existence of an optimal quantum speedup logically implies the mutually exclusive $or$ of well defined causal loops. Of course, these causal loop must appear in the quantum description of the quantum algorithm that delivers the quantum speedup that, involving only the measurements of commuting observables, is submitted to the same (classical) logic. In turn, it is easily seen that the only way of obtaining these causal loops is by time-symmetrizing the ordinary quantum description of the quantum algorithm. In fact, to have a causal loop, we must properly share the selection of the pair problem-setting and corresponding solution between the initial and final measurements. Because of the time-symmetry of the unitary evolution between the two one-to-one correlated measurement outcomes, we should share it evenly. This exactly specifies the time-symmetrization we are dealing with. Therefore, as long as logic and quantum mechanics are consistent, the time-symmetric quantum causal loops generated by the time-symmetrization operation should be exempt from the paradoxes that plague the usual kind.
Interestingly, the time-symmetrization we are dealing with turns out to be mandatory also to satisfy a symmetry implicit in any unitary evolution between two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement.
@@
In fact, in this kind of process, we can postpone the initial measurement at will – all results must remain unaltered. When the two measurements become simultaneous, the time-symmetrization in question must already be there to become the atemporal symmetrization mandatory to represent the perfect symmetry between the simultaneous measurements of two maximally entangled observables.
This further motivation for time-symmetrizing the unitary evolutions we are dealing with, besides further justifying the previous motivation, makes things much more general. Any such unitary evolution must be time-symmetrized, what leaves it mathematically unaltered but changes its ordinary, causal, quantum description into that a quantum superposition of time-symmetric causal loops.
If the unitary evolution is a quantum algorithm that determines the solution with certainty, in each causal loop it is as if the problem solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future. However, this just creates the room for its occurring with quantum speedup; the quantum algorithm does not need any more to be optimal like in the case of Grover algorithm. How much of this room is exploited depends on the problem-solver, i.e. on the quantum algorithm.
If the unitary evolution in question is not a quantum algorithm, but anyhow creates one-to-one quantum correlation between two initially unrelated observables (without changing the eigenvalue of the observable selected by the initial measurement), it can always occur with quantum speedup. In this more general case, the problem-solver advanced knowledge might loose any meaning but, of course, can be replaced by the reduction of the dimension of the Hilbert space in which the search for one-to-one correlation takes place. This reduction, in turn, creates the room for a quantum speedup.
There is also the case that the unitary evolution does not build up any quantum correlation. It is the case of quantum nonlocality, namely the measurement in the same basis of two maximally entangled observables whose respective subsystems are in a state of spatial separation. Here maximal entanglement between the two observables is already there before the space-separation process. After the time-symmetrization, this process becomes a quantum superposition of time-symmetric quantum causal loops locally going from one to the other measurement outcome via the time the two subsystems were not yet spatially separate. Here, always locally, these causal loops tell either observable the state to assume in view of its future measurement – see $\left[5\right]$. We can see that these causal loops play the role of the hidden variables envisaged by Einstein and the others $\left[11\right]$ to locally (i.e. also physically) describe quantum nonlocality.
Note that, as a consequence, the incompleteness we are currently ascribing to the quantum description of quantum algorithms is the same that Einstein and the others ascribed to the quantum description of quantum nonlocality. In both cases, what is incomplete is the quantum description of a quantum correlation. With this, we end the first part of the work.
2.2 Teleological character of natural evolutions – a possible physical basis

### v6

@@ 1 Introduction
In this work we propose a physical basis for the notion of teleological evolution, that is, an evolution toward a goal with an attractor in the same goal it will produce in the future. It is, of course, a puzzling notion since it implies a causal loop. For centuries it has been applied by natural philosophy to the evolutions of the living; it was then dismissed with the advent of modern science because of its alleged lack of a physical basis.
What we think should change that (also centuries old) position of science is the relatively recent discovery of the quantum computational speedup. In this work, by quantum (computational) speedup we mean its strict form, that is, the fact that quantum algorithms solve problems with fewer logical operations (computation steps) than classically possible.
@@
As we try to be interdisciplinary, let us provide an example of quantum speedup right away. Consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem. If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case – by the optimal quantum search algorithm of Grover $\left[1\right]$ – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of seeing the thing.
@@
Let us mention priorities along the path that led to this fundamental discovery. The notion that computation is possible in the quantum world, with the related notion of quantum bit, was introduced by Finkelstein $[2]$ in 1969. In 1982, Feynman $\left[3\right]$ highlighted the higher computational efficiency of the quantum world, showing that the classical simulation of a quantum process can require an amount of physical resources exponentially higher than the amount involved in the quantum process itself. In 1985, Deutsch $\left[4\right]$ devised the seminal quantum algorithm that delivers a quantum computational speedup.
The reason we believe that that position of science should be changed is simply the fact that quantum algorithms turn out to be teleological evolutions, that is, evolutions toward a goal (the solution of the problem) with an attractor in the same goal they will produce in the future (the solution again). We will show that this is the immediate consequence of a retrocausal approach to the quantum speedup of ours and co-authors $\left[5-9\right]$.
The present work is divided into two parts. The first is a review and further clarification of that retrocausal approach to the quantum speedup: it will justify the teleological character of quantum algorithms. The second is whether the same justification could constitute the physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived. The present introduction is divided into two corresponding parts.
1.1 The teleological character of quantum algorithms
Let us focus on Grover quantum search algorithm, by the way a most universal quantum algorithm that can solve with a quadratic quantum speedup any NP problem.
We show that just opening to retroausality allows to produce a simple quantitative explanation of the optimal quantum speedup of Grover algorithm that will also apply to a very large class of quantum speedups.
Let us begin with the four drawers case. Our argument goes as follows.
@@ 1.1 The teleological character of quantum algorithms
The fact that Alice always solves the problem of locating the ball by opening just one drawer is logically equivalent to her knowing in advance, before she begins her problem-solving action, that the ball is in a given pair of drawers and benefiting of that knowledge to solve the problem by opening just one drawer. In fact, if the ball is in the drawer opened, she has solved the problem; if not, it must be in the other drawer and she has solved the problem of locating the ball as well, in either case having opened just one drawer. Of course, what Alice should know in advance is the mutually exclusive $or$ of the three possible ways of pairing the number of the drawer with the ball with that of another drawer. Note that this also means knowing in advance half of the information that specifies the solution of the problem.
@@
Then the question becomes: from where does Alice, before beginning her problem-solving action, always get that information about the solution? Since, to her, the only possible source of it is the solution of the problem she will produce and measure in the future, the only possible answer is that it comes to her from her future measurement of the solution. Of course, this implies retrocausality, more specifically a causal loop.
@@
The above immediately generalizes to any number, $N$, of drawers. Grover algorithm yields a quadratic quantum speedup, that is, always solves the problem by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers against the $\operatorname*{O}\left(N\right)$ of the best classical case. Also now, this quantum speedup can be logically justified by the fact that Alice knows in advance, in a mutually exclusive way, one of the possible halves of the information about the number of the drawer with the ball. This means that she knows in advance that the ball is (in a mutually exclusive way) in one of the possible tuples of $\sqrt{N}$ drawers. She then solves the problem in the best possible logical way by opening $\operatorname*{O}\left(\sqrt{N}\right)$ drawers, which is indeed what Grover algorithm does.
We can conclude that the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of well defined causal loops, each for each possible way of taking half of the information about the solution. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution by opening fewer drawers.
We can formulate the following advanced knowledge rule that, as checked in $\left[8\right]$, also holds for all the major quantum algorithms for the solution of oracle problems:
@@
The number of computation steps (oracle queries) needed to solve the (oracle) problem in an optimal quantum way is the number needed to solve it in an optimal logical way benefiting of the advanced knowledge of one (any one) of the possible halves of the information about the solution.
Later we will show why this rule applies to any problem for which a problem setting and a corresponding solution can be identified. By the way, it is a rule that does not exist in mainstream quantum computing science. There, the only way of knowing the number of computation steps needed to solve a problem with a quantum speedup is finding a quantum algorithm that does that. As is well known, this is a matter of high ingenuity. Of course, this goes along with the fact that mainstream quantum computing science, as mainstream quantum mechanics, relies on the notion of causal evolution that is incompatible with that of causal loop.
Let us make a digression. As the work goes on, we will naturally go into more and more detail, with the consequent risk of masking the forest with the trees. So, we would like to emphasize that the forest is right now in front of our eyes. The notion of causal loops is obviously controversial; probably most people believe that causal loops cannot exist except perhaps in very exotic physical situations such as the closed timelike curves allowed by general relativity. Now we would have, moreover trivially, demonstrated that the quantum speedup of a quantum algorithm logically implies their existence. This is the forest! As we will venture into very exotic situations ourselves, but in ordinary non-relativistic quantum mechanics, we would like to suggest that we keep this basic fact in mind: it is the source of all exoticism.
Back to our line of thinking, of course we should ask ourselves what is the impact of the fact that a quantum speed up logically implies the mutually exclusive $or$ of causal loops on our way of seeing the quantum algorithm, that is, on the quantum description that we make of it.
To this end, we should make reference to a quantum description of the quantum algorithm less incomplete than the usual one, which is limited to the process of solving the problem (where the quantum speedup resides) and lacks the initial measurement. Of course, the canonical quantum description of a quantum process must include the initial and final measurements and the unitary evolution in between.
Of course, this problem is easily adjusted. It suffices to extend the usual quantum description of the process of solving the problem to the upstream process of setting the problem. The extended quantum description is thus as follows. An initial measurement, in a quantum superposition (or mixture) of all the possible problem settings, selects one of them at random222 The problem-setter would be free to unitarily change the randomly selected problem-setting into a wanted one. We disregard this possibility that, in the present context, changes nothing.      . This is followed by the unitary computation of the corresponding solution and by the final measurement that reads it without changing the quantum state333 With four drawers, Grover algorithm yields the solution of the problem with certainty; when the drawers are more than four, by Grover algorithm we will always mean Long’s version of it $\left[10\right]$, which always provides the solution with certainty. . Note that there is one-to-one correlation between the initial measurement outcome, which encodes the problem-setting, and the final one, which encodes the corresponding solution – in Grover algorithm, they are both the number of the drawer with the ball. Furthermore, there is a unitary evolution connecting the two measurement outcomes, despite the fact it comprises the final measurement (which indeed does not change the quantum state). This extension of the usual quantum description of quantum algorithms to the process of setting the problem will play a key role in the present work. We will call it the ordinary quantum description of the quantum algorithm, ordinary because it is still the ordinary causal description of mainstream quantum mechanics.
We can go now to the impact of the logical implication in question on our way of describing the quantum algorithm.
The first thing to observe is that, involving only the measurements of commuting observables (the problem-setting and the solution observables), the quantum algorithm is subject to the same (classical) logic of the logical implication in question; in other words, the non-classical quantum logic of non-commuting observables is not involved here. The quantum algorithm that delivers the optimal quantum speedup must therefore accommodate the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its quantum speedup (in the quantum world, the logical mutual exclusive $or$ of possibilities becomes their quantum superposition). However, being causal in character, its ordinary quantum description cannot describe causal loops. According to the present line of thinking, we must conclude that the ordinary quantum description of the quantum algorithm is either wrong or incomplete, since it does not and, being causal, cannot describe causal loops. We will see that it is incomplete as it is completed by time-symmetrizing it.
We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes – the problem-setting and the corresponding solution. The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement, that performed by the final measurement, time-symmetrically, should propagate backward in time until replacing the previous initial measurement outcome. This is a causal loop that changes its own past; it is, indeed, one of the causal loops logically implied by the sheer existence of the quantum speedup. Of course, we should then take the quantum superposition of all the possible ways of evenly sharing. We will see that all this, mathematically, leaves the unitary part of the ordinary quantum description of the quantum algorithm unaltered but changes the description of the behavior of causality along that part. The single, unitary, causal process of the ordinary quantum description of the quantum algorithm changes into the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
We will (easily) show that this time-symmetrization is the only possible way to generate the quantum superposition of the causal loops in question, which in turn must be hosted within the quantum algorithm. Therefore, according to the current line of thinking, it is imposed by logic and quantum mechanics.
Most importantly to present ends, the fact that, in each causal loop, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution with fewer computation steps highlights the teleological character of quantum algorithms. They are indeed evolutions toward a goal (the solution) with an attractor in the very goal they will produce in the future (the solution again).
Until now we have focused on Grover quantum search algorithm, but the results obtained go well beyond it.
In the first place, we never needed to know the unitary evolution between the two one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement (i.e., Grover algorithm). It sufficed to know that such an evolution exists.
In the second place, until now, the time-symmetrization of the ordinary quantum description of the quantum algorithm has been imposed by the necessity of generating the quantum superposition of the causal loops whose mutually exclusive $or$ is logically implied by the sheer existence of its quantum speedup. However, it turns out that there is another reason for time-symmetrizing, independent of this.
@@
Let us consider the unitary (thus reversible) part of the quantum algorithm. It is a unitary evolution interconnecting two one-to-one correlated measurement outcomes that never changes the eigenvalue selected by the initial measurement. This allows to postpone the initial measurement at will. When made simultaneous with the final measurement, the time-symmetrization in question must already be in place to become the perfect atemporal symmetry between the simultaneous measurements of two maximally entangled observables.
The two things together make the results obtained in the case of Grover algorithm applicable to a very general class of quantum processes involving quantum correlation. In fact, they imply that:
Any unitary evolution between an initial and final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing its ordinary quantum description.
We call the above the causal loop rule of quantum correlation.
We can see that it is a generalization also when the process is still a quantum algorithm. In fact, until now we have remained committed to Grover algorithm that is an optimal one. However, also the time-symmetrization of the ordinary quantum description of non-optimal quantum algorithms yields a quantum superposition of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution. The point is that this time she may not exploit this information at all, or not fully exploit it. For example, the latter is the case of the random quantum search algorithms $\left[11\right]$, which provides a suboptimal quantum speedup. We will be interested in this kind of quantum algorithms further on.
In view of applying the causal loop rule of quantum correlation to processes that have the form of quantum algorithms but are not that, let us note that we can do without the anthropomorphic notion of a problem-solver that produces the solution with fewer computation steps because she knows in advance half of the information about it. We can also avoid to speak of a “quantum algorithm”. We can say that, in each time-symmetrization instance, there is a reduction of the dimension of the Hilbert space in which the search for the one-to-one correlation between the initial and final measurement outcomes takes place and consequently the correlation in question can be achieved with a quantum speedup.
Of course, the process does not even need to have the form of a quantum algorithm. In other words, it does not need to create one-to-one correlation between two initially unrelated observables. The two observables can be maximally entangled before the process begins. As we shall see, quantum nonlocality is exactly a process of this kind and the quantum superposition of causal loops generated by its time-symmetrization explains the EPR (Einstein, Podolsky, and Rosen) paradox.
1.2 The possible teleological character of the evolutions of the living